# Cross-Comparison of Sampling Algorithms for Pulse Profile Modeling of PSR J0740+6620

### v1

@@ I Introduction
Neutron stars are among the densest objects in the Universe, making them excellent for studying the behavior of dense nuclear matter and constraining the equation of state (EoS) (see e.g., Lattimer and Prakash, 2016; Baym et al., 2018; Tolos and Fabbietti, 2020). One effective approach to probing NS properties, and thereby constraining the EoS, is through modeling the X-ray emission from their surfaces. The Neutron Star Interior Composition Explorer (NICER; Gendreau et al. (2016)), located onboard the International Space Station, has been pivotal in this effort by detecting soft thermal X-rays from rotation-powered millisecond pulsars. These pulsars are a distinct category of neutron stars, showing rotationally modulated and hence pulsed X-ray emission that is thought to originate from heat deposited at the magnetic poles by return currents (see e.g., Ruderman and Sutherland, 1975; Arons, 1981; Harding and Muslimov, 2001). By analyzing NICER observations using pulse profile modeling (e.g., Watts, 2019; Bogdanov et al., 2019, 2021, and references therein) — a technique that incorporates relativistic effects caused by the neutron star’s rapid spin and strong gravitational field — it is possible to derive precise measurements of neutron star masses and radii. These measurements, in turn, provide stringent constraints on the EoS governing cold, ultra-dense matter (see e.g. Rutherford et al., 2024; Koehn et al., 2024; Huang et al., 2025; Golomb et al., 2025; Li et al., 2024).
@@ II.1 Pulse Profile Modeling using X-PSI
We use the open-source X-PSI package, with versions ranging from v2.2.0 to v2.2.7 Riley et al. (2023). As done previously (see e.g., Riley et al., 2021; Salmi et al., 2022, 2024a), we use the ‘Oblate Schwarzschild + Doppler’ approximation to model the energy-resolved X-ray pulses from the NS Miller and Lamb (1998); Nath et al. (2002); Poutanen and Gierliński (2003); Morsink et al. (2007); Lo et al. (2013); AlGendy and Morsink (2014); Bogdanov et al. (2019); Watts (2019). This approximation takes the oblate shape of the NS into account and the relativistic effects resulting from the rapid spin of the rotation-powered millisecond pulsars while treating the exterior spacetime as Schwarzschild. See Riley et al. (2019, 2023) for a more detailed description of X-PSI.
For PSR J0740+6620, we use the priors for mass, inclination, and distance from Fonseca et al. (2021). Additionally, we use the TBabs model (Wilms et al. (2000), updated in 2016) to account for the absorption of X-rays in the interstellar medium. We use the fully ionized hydrogen atmosphere model NSX Ho and Lai (2001); Ho and Heinke (2009), an assumption which does not seem to have a significant effect on the inferred radius of PSR J0740+6620 Salmi et al. (2023). Furthermore, the hot emitting regions of the NS are modeled with two circular uniform temperature regions using the ST-U (Single-Temperature-Unshared) model (see Vinciguerra et al. (2023); Riley et al. (2019) for more details on the X-PSI model naming convention and parameters). To compute the likelihood we adopt either high or low-resolution settings 222 Low-resolution settings: num_leaves=32, num_energies=32, sqrt_num_cells=18, and max_sqrt_num_cells=32. High-resolution settings: num_leaves=64, num_energies=128, sqrt_num_cells=32, and max_sqrt_num_cells=64. for our inference runs (for details see see Section 2.3.1 of Vinciguerra et al. (2023)). These settings control the discretization of the computational domain for computation of signals (pulse profiles) incident on the telescope. In this study, we mostly opt for low-resolution settings, because they significantly reduce computational cost while having little to no noticeable impact on the resulting posteriors.
@@ III.1 Real Data of PSR J0740+6620
We use the NICER data of PSR J0740+6620 collected between September 21, 2018, and April 21, 2022 (observation IDs 1031020101 through 5031020445). The NICER X-ray event data is processed in the exact same manner as reported in Salmi et al. (2024a). After filtering, the final dataset includes 2.73381 Ms of on-source exposure time, representing an increase of more than 1 Ms compared to the “old” NICER dataset (2018 September 21 -- 2020 April 17) used in Riley et al. (2021); Miller et al. (2021).
@@ V Discussion & Future Work
In this study, we used the X-PSI framework to assess the robustness of mass and radius inferences concerning our choice of sampler for the high-mass pulsar PSR J0740+6620, building on the work of Salmi et al. (2024a). We focused on two sampling algorithms: the multimodal nested sampling algorithm implemented in MultiNest (used in Salmi et al. (2024a)) and the slice nested sampling algorithm implemented in UltraNest. Both samplers performed reliably, producing accurate and unbiased parameter estimation results when analyzing simulated data. Additionally, both samplers provided consistent results for the joint NICER and XMM-Newton PSR J0740+6620 data, with only slight deviations in the posterior tails (see Figure 4). This consistency shows that the choice of sampler does not significantly affect the inferred properties of PSR J0740+6620 and that the results are robust.
@@ Acknowledgements.
M.H. and A.L.W. acknowledge support from the NWO grant ENW-XL OCENW.XL21.XL21.038 Probing the phase diagram of Quantum Chromodynamics (PI: Watts). T.S., Y.K., and A.L.W. acknowledge support from ERC Consolidator grant No. 865768 AEONS (PI: Watts). The NWO Domain Science subsidized the use of the national computer facilities in this research. Part of the work was also carried out on the HELIOS cluster including dedicated nodes funded via the above-mentioned ERC Consolidator grant. This research has made use of data products and software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory. Software: Cython (Behnel et al., 2011), GetDist (Lewis, 2019), GNU Scientific Library (Galassi et al., 2009), HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc), Matplotlib (Hunter, 2007), MPI for Python (Dalcín et al., 2008), MultiNest (Feroz et al., 2009), nestcheck (Higson, 2018), NumPy (van der Walt et al., 2011), PyMultiNest (Buchner, 2016b), Python/C language (Oliphant, 2007), SciPy (Jones et al., 2001), UltraNest (Buchner, 2021), and X-PSI (Riley et al., 2023). Facilities: NICER, XMM-Newton

### v2

@@ I Introduction
Neutron stars are among the densest objects in the Universe, making them excellent for studying the behavior of dense nuclear matter and constraining the equation of state (EoS) (see e.g., Lattimer and Prakash, 2016; Baym et al., 2018; Tolos and Fabbietti, 2020). One effective approach to probing neutron star properties, and thereby constraining the EoS, is through modeling the X-ray emission from their surfaces. The Neutron Star Interior Composition Explorer (NICER; Gendreau et al. (2016)), located onboard the International Space Station, has been pivotal in this effort by detecting soft thermal X-rays from rotation-powered millisecond pulsars. These pulsars are a distinct category of neutron stars, showing rotationally modulated and hence pulsed X-ray emission that is thought to originate from heat deposited at the magnetic poles by return currents (see e.g., Ruderman and Sutherland, 1975; Arons, 1981; Harding and Muslimov, 2001). By analyzing NICER observations using pulse profile modeling (e.g., Watts, 2019; Bogdanov et al., 2019, 2021, and references therein) — a technique that incorporates relativistic effects caused by the neutron star’s rapid spin and strong gravitational field — it is possible to derive precise measurements of neutron star masses and radii. These measurements, in turn, provide stringent constraints on the EoS governing cold, ultra-dense matter (see e.g. Rutherford et al., 2024; Koehn et al., 2024; Huang et al., 2025; Golomb et al., 2025; Li et al., 2024).
@@ II.1 Pulse Profile Modeling using X-PSI
We use the open-source X-PSI package, with versions ranging from v2.2.0 to v2.2.7 Riley et al. (2023). As done previously (see e.g., Riley et al., 2021; Salmi et al., 2022, 2024a), we use the ‘Oblate Schwarzschild + Doppler’ approximation to model the energy-resolved X-ray pulses from the neutron star Miller and Lamb (1998); Nath et al. (2002); Poutanen and Gierliński (2003); Morsink et al. (2007); Lo et al. (2013); AlGendy and Morsink (2014); Bogdanov et al. (2019); Watts (2019). This approximation takes the oblate shape of the neutron star into account and the relativistic effects resulting from the rapid spin of the rotation-powered millisecond pulsars while treating the exterior spacetime as Schwarzschild. See Riley et al. (2019, 2023) for a more detailed description of X-PSI.
For PSR J0740+6620, we use the priors for mass, inclination, and distance from Fonseca et al. (2021). Additionally, we use the TBabs model (Wilms et al. (2000), updated in 2016) to account for the absorption of X-rays in the interstellar medium. We use the fully ionized hydrogen atmosphere model NSX Ho and Lai (2001); Ho and Heinke (2009), an assumption which does not seem to have a significant effect on the inferred radius of PSR J0740+6620 Salmi et al. (2023). Furthermore, the hot emitting regions of the neutron star are modeled with two circular uniform temperature regions using the ST-U (Single-Temperature-Unshared) model (see Vinciguerra et al. (2023); Riley et al. (2019) for more details on the X-PSI model naming convention and parameters). To compute the likelihood we adopt either high or low-resolution settings 222 Low-resolution settings: num_leaves=32, num_energies=32, sqrt_num_cells=18, and max_sqrt_num_cells=32. High-resolution settings: num_leaves=64, num_energies=128, sqrt_num_cells=32, and max_sqrt_num_cells=64. for our inference runs (for details see see Section 2.3.1 of Vinciguerra et al. (2023)). These settings control the discretization of the computational domain for computation of signals (pulse profiles) incident on the telescope. In this study, we mostly opt for low-resolution settings, because they significantly reduce computational cost while having little to no noticeable impact on the resulting posteriors.
@@ III.1 Real Data of PSR J0740+6620
We use the NICER data of PSR J0740+6620 collected between September 21, 2018, and April 21, 2022 (observation IDs 1031020101 through 5031020445). The NICER X-ray event data is processed in the exact same manner as reported in Salmi et al. (2024a). After filtering, the final dataset includes 2.73 Ms of on-source exposure time, representing an increase of more than 1 Ms compared to the “old” NICER dataset (2018 September 21 -- 2020 April 17) used in Riley et al. (2021); Miller et al. (2021).
@@ V Discussion & Future Work
In this study, we used the X-PSI framework to assess the robustness of mass and radius inferences concerning our choice of sampler for the high-mass pulsar PSR J0740+6620, building on the work of Salmi et al. (2024a). We focused on two sampling algorithms: the multimodal nested sampling algorithm implemented in MultiNest (used in Salmi et al. (2024a)) and the slice nested sampling algorithm implemented in UltraNest. Both samplers performed reliably, producing accurate and unbiased parameter estimation results when analyzing simulated data. Additionally, both samplers provided consistent results for the joint NICER and XMM-Newton PSR J0740+6620 data, with only slight deviations in the posterior tails (see Figure 4). This consistency shows that the choice of sampler does not significantly affect the inferred properties of PSR J0740+6620.
@@ Acknowledgements.
M.H. and A.L.W. acknowledge support from the NWO grant ENW-XL OCENW.XL21.XL21.038 Probing the phase diagram of Quantum Chromodynamics (PI: Watts). Y.K., T.S., and A.L.W. acknowledge support from ERC Consolidator grant No. 865768 AEONS (PI: Watts). The NWO Domain Science subsidized the use of the national computer facilities in this research. Part of the work was also carried out on the HELIOS cluster including dedicated nodes funded via the above-mentioned ERC Consolidator grant. This research has made use of data products and software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory. Software: Cython (Behnel et al., 2011), GetDist (Lewis, 2019), GNU Scientific Library (Galassi et al., 2009), HEASoft (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc), Matplotlib (Hunter, 2007), MPI for Python (Dalcín et al., 2008), MultiNest (Feroz et al., 2009), nestcheck (Higson, 2018), NumPy (van der Walt et al., 2011), PyMultiNest (Buchner, 2016b), Python/C language (Oliphant, 2007), SciPy (Jones et al., 2001), UltraNest (Buchner, 2021), and X-PSI (Riley et al., 2023). Facilities: NICER, XMM-Newton